# Style-NeRF2NeRF: 3D Style Transfer from Style-Aligned Multi-View Images

### v2

@@ 4 Results
We run our experiments on several real-world scenes, including the Instruct-NeRF2NeRF Haque et al. (2023) dataset captured by a smartphone and a mirrorless camera with camera poses extracted by COLMAP Schönberger and Frahm (2016) and PolyCam pol . The dataset contains large-scale 360 scenes, objects, and forward-facing human portraits. We show qualitative results and comparisons against several variants to verify the effectiveness of our method design with CLIP Text-Image Direction Similarity (CLIP-TIDS), a quantitative metric originally introduced in StyleGAN-Nada Gal et al. (2022) and CLIP Directional Consistency (CLIP-DC), a score proposed by Instruct-NeRF2NeRF Haque et al. (2023) that aims to measure the view-consistency between adjacent frames. In addition to the above, we also present comparison results against recent NeRF-based 3D editing methods, Instruct-NeRF2NeRF Haque et al. (2023) and ViCA-NeRF Dong and Wang (2024). We encourage our readers to see the results in the supplementary video.
@@
Figure 5: Baseline Comparisons: We compare our method against several variants. The images show an example comparison of the "Bear" scene edited with a text description "A bear, impressionism painting style" with a text guidance scale of $7.5$. Note that (b), (c), (e), and (f) are all novel view renders from NeRF. We encourage our readers to check the results in the video.
@@
Figure 6: Qualitative Results: We show novel view rendering examples of real-world scenes stylized or edited with text descriptions specifying certain artistic styles or environmental changes such as weather conditions.
@@
Figure 7: Method Comparison: A comparison of NeRF stylization methods. While we used a text guidance scale of between $15$ to $25$ for our results, it is controllable via text prompts concerning subjective preferences. Note that all images are novel view renders from NeRF.
4.1 Qualitative Evaluation
@@ 4.2 Ablations
4.2.1 Quantitative Evaluation.
We quantitatively measure our method against the variants using CLIP-TIDS and CLIP-DC with a fixed text guidance scale of 15. The results are shown in table 1.
Table 1: Quantitative Evaluation We show both CLIP-TIDS and CLIP-DC results. The values are the average of the novel view render frames over two scenes using five prompts.
@@
CLIP-TIDS $0.125$ $0.073$ $0.237$ $\mathbf{0.162}$
CLIP-DC 0.932 0.917 0.917 0.928
4.3 Method Comparison
We compare our method against Instruct-NeRF2NeRF Haque et al. (2023) and ViCA-NeRF Dong and Wang (2024) on four scenes including two large-scale outdoor scenes, a 360 object scene (Bear) and a forward-facing scene (a human portrait) using three different text prompts for each scene. Our method exhibits competitive style transfer results whereas previous methods occasionally suffer from hallucination effects (e.g. The Janus problem etc…) caused by the underlying diffusion model. As the generation of images and NeRF refinement is a separate process in our method, it is possible to filter out and recreate any images that could have undesired impact on the NeRF fine-tuning. Visual results are given in figure 7.
4.3.1 User Study
@@ 4.3.2 Quantitative Comparison
Table 2: Method Comparison Results The metrics are the average of novel view renders over four scenes with each using three prompts. (Average of $4\times 3=12$ style transfer results.) Our method shows the best values for CLIP-TIDS, CLIP-DC, and user preference.
CLIP-TIDS CLIP-DC User Preference
Instruct-NeRF2NeRF $0.081$ $0871$ $15.3\%$
@@
5 Limitations and Future Work
While our method may apply artistic style transfer to various scenes, including large-scale outdoor environments, there are several limitations that we’d like to consider. Thin structures such as plants and trees in the background or delicate texture patterns are challenging to reconstruct due to ambiguities in the stylized multi-view images. For the same reason, our method will struggle to learn fine details if there is too much variation in the training images (e.g. different people or objects in the background, random patterns of clouds in the sky). As the style-aligned diffusion pipeline is conditioned on depth maps, significant editing of geometry is also difficult.
We think our approach is applicable to other types of 3D representations such as 3D Gaussian Splatting Kerbl et al. (2023) and extendable to more features such as scene relighting and deformation, which are exciting directions for further exploration.
6 Conclusion
@@
7 Acknowledgements
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015, and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. We want to thank Instruct-NeRF2NeRF Haque et al. (2023) authors for sharing their dataset, and Yuki Kato, Shinji Terakawa, and Yoshiaki Tahara for assisting with data capturing.
@@ Appendix A Additional Implementation Details
We pre-train the "nerfacto" model implemented in Nerfstudio Tancik et al. [2023] for 60,000 iterations and then fine-tune for 15,000 iterations. We use the default "nerfacto" losses; RGB pixel loss $\mathcal{L}_{rgb}$, distortion loss $\mathcal{L}_{dist}$ Barron et al. [2022], interlevel loss $\mathcal{L}_{inter}$ Barron et al. [2022], orientation loss $\mathcal{L}_{orien}$ Verbin et al. [2022], and predicted normal loss $\mathcal{L}_{normal}$ Verbin et al. [2022] for pre-training (equation 12). During fine-tuning, we disable the RGB pixel loss $\mathcal{L}_{rgb}$, orientation loss $\mathcal{L}_{orien}$, and predicted normal loss $\mathcal{L}_{normal}$ but add the Sliced Wasserstein Distance (SWD) loss Heitz et al. [2021] (equation 13). The total loss function for each phase is as follows:
$\mathcal{L}_{pre}=\mathcal{L}_{rgb}+\lambda_{1}\mathcal{L}_{dist}+\lambda_{2}\mathcal{L}_{inter}+\lambda_{3}\mathcal{L}_{orien}+\lambda_{4}\mathcal{L}_{normal}$ (12)
$\mathcal{L}_{fine}=\mathcal{L}_{swd}+\lambda_{1}\mathcal{L}_{dist}+\lambda_{2}\mathcal{L}_{inter}$ (13)
where we use the default hyper-parameters $\lambda_{1}=0.002,\lambda_{2}=1.0,\lambda_{3}=0.0001,\lambda_{4}=0.001$ for most cases. A greater value for the distortion loss may work better if floating artifacts remain in the scene. We list a brief description of the "nerfacto" losses.
  • Distortion Loss: The loss encourages the density along a ray to become compact, aiming to prevent floaters and background collapse. It was proposed in Mip-NeRF 360 Barron et al. [2022].
  • Interlevel Loss: The loss allows the histograms of the point sampling proposal network and NeRF network to become more consistent. It was also proposed in Mip-NeRF 360 Barron et al. [2022]
  • Orientation Loss: The loss aims to prevent "foggy" surfaces by penalizing visible samples with predicted normals facing the ray direction. It was introduced in Ref-NeRF Verbin et al. [2022].
@@
For detailed definitions of the "nerfacto" losses, please see Mip-NeRF Barron et al. [2022] and Ref-NeRF Verbin et al. [2022].
The SWD loss $\mathcal{L}_{swd}$ is applied to $64\times 64$ patches sampled during NeRF fine-tuning. Although we found that $64\times 64$ empirically works sufficiently, one may change the patch size accordingly. We run all our experiments with Python 3.10 and CUDA 11.8 on a single NVIDIA H100. Although we optionally enable SDEdit Meng et al. [2021] in our style-aligned diffusion pipeline, we recognized that depth maps are often enough for conditioning on the original views. In such cases, we use $1.0$ as the strength for SDEdit.
Appendix B CLIP Text-Image Direction Similarity
CLIP Text-Image Direction Similarity (CLIP TIDS) Gal et al. [2022] is a metric for evaluating how well the change in the stylized image is aligned with the user-provided text prompt. Given the CLIP image and text encoder $E_{I},E_{T}$, CLIP TIDS between the source and the stylized image $I_{source},I_{style}$ is calculated as:
$\Delta I=E_{I}(I_{style})-E_{I}(I_{source}),\Delta T=E_{T}(t_{style})-E_{T}(t_{source})$ (14)
@@ Appendix B CLIP Text-Image Direction Similarity
where $t_{source},t_{style}$ are text descriptions describing the images (e.g. $t_{source}=$"A photo of a person", $t_{style}=$"A person like Vincent Van Gogh").
Appendix C Effect of Text Guidance Scale
In figure 8, we show an example of NeRF renderings using the same style prompt "A person like Vincent Van Gogh" but different text guidance scales. We may see that style strength is controllable while keeping the original content structure grounded on the original image. We also show CLIP TIDS for each text guidance scale in table 3. A stronger text guidance scale will lead to higher CLIP TIDS.
Table 3: CLIP-TIDS Comparison CLIP-TIDS values over test renders for each text guidance scale are shown. We can verify that a stronger text guidance scale results in higher CLIP TIDS values.
Text Guidance Scale $s=7.5$ $s=15.0$ $s=22.5$ $s=30.0$ $s=37.5$
@@ Appendix C Effect of Text Guidance Scale
Figure 8: Effect of Text Guidance Scale: We show some NeRF rendering results for the prompt "A person like Vincent Van Gogh" using various text guidance scales ($s=7.5,15.0,22.5,30.0,37.5$).
Appendix D Limitations
While our method can effectively perform overall style transfer to 3D scenes, it is still difficult to reconstruct a detailed structure of fluctuating objects within the stylized multi-view images. In figure 9, for example, (c) the clouds in the fine-tuned NeRF scene show a fractal-like pattern, which is different from (b) the clouds illustrated in the stylized image. This phenomenon is due to the ambiguities of cloud positions or shapes appearing in the stylized images generated by the style-aligned diffusion pipeline Hertz et al. [2023b]. We leave the development of a more robust content structure-preserving style transfer technique as future work.
Figure 9: Limitations: Due to remaining ambiguities in the stylized multi-view images, fluctuating objects such as clouds may lose their detailed shape in the fine-tuned NeRF renderings. We wish to improve on this in our future work.

### v3

@@ 4. Results
We show qualitative results and comparisons against several variants to verify the effectiveness of our method design with CLIP Text-Image Direction Similarity (CLIP-TIDS), a metric introduced initially in StyleGAN-Nada (Gal et al., 2022) and CLIP Directional Consistency (CLIP-DC), a score proposed by Instruct-NeRF2NeRF (Haque et al., 2023) that aims to measure the directional similarity between original and stylized views. We also evaluate the temporal view consistency (Lai et al., 2018) of the stylized 3D scenes by calculating the average warping error between adjacent frames using FlowNet2 (Ilg et al., 2017), an off-the-shelf optical flow estimation model. In addition to the above, we present comparison results against recent NeRF-based 3D editing methods, Instruct-NeRF2NeRF (Haque et al., 2023) and ViCA-NeRF (Dong and Wang, 2024). We encourage our readers to see the results in the supplementary video.
@@
Figure 6. Baseline Comparisons. We compare our method against several variants. The images show an example comparison of the ”Bear” scene trained from a style description ”A water painting of a brown bear” with a text guidance scale of $7.5$. Note that (b), (c), (e), and (f) are all novel view renders from NeRF. NeRF renderings from (f) ours preserve the original content in (a) without noticeable artifacts compared to (c) Train-from-Scratch and (e) Style-Alignment w/RGB Loss, and also maintain style and color similar to the 2D reference (d). Unlike ours, No Style-Alignment (b) fails to preserve consistent scene color. We encourage our readers to check the results in the video.
@@
Figure 7. Qualitative Results. We show novel view rendering examples of real-world scenes stylized or edited with text descriptions specifying certain artistic styles or environmental changes such as weather conditions.
@@
Figure 8. Method Comparison. A comparison of NeRF stylization methods. While we used a text guidance scale of between $15$ to $25$ for our results, it is controllable via text prompts concerning subjective preferences. Note that all images are novel view renders from NeRF.
4.1. Qualitative Evaluation
@@ 4.2. Ablations
4.2.1. Quantitative Evaluation.
We quantitatively measure our method against the variants using CLIP-TIDS, CLIP-DC, and the average warping error with a fixed text guidance scale of 15. The results are shown in table 1.
Table 1. Quantitative Evaluation. We show CLIP-TIDS, CLIP-DC, and the averaged warping error (MSE) measured across rendered view frames from novel camera trajectories. The values are the average of two scenes using five prompts.
@@
CLIP-TIDS $\uparrow$ $0.125$ $0.073$ $0.127$ $\mathbf{0.162}$
CLIP-DC $\uparrow$ $\mathbf{0.932}$ 0.917 0.917 0.928
Warp Error $\downarrow$ 0.345 0.367 0.351 $\mathbf{0.337}$
4.3. Method Comparison
We compare our method against Instruct-NeRF2NeRF (Haque et al., 2023) and ViCA-NeRF (Dong and Wang, 2024) on four scenes including two large-scale outdoor scenes, a 360 object scene (Bear) and a forward-facing scene (a human portrait) using three different text prompts for each scene. Our method exhibits competitive style transfer results whereas previous methods occasionally suffer from hallucination effects (e.g. The Janus problem etc…) caused by the underlying diffusion model. As the generation of images and NeRF refinement is a separate process in our method, it is possible to filter out and recreate any images that could have undesired impact on the NeRF fine-tuning. Visual results are given in figure 8.
4.3.1. User Study
@@ 4.3.2. Quantitative Comparison
Table 2. Method Comparison Results. The metrics are the average of novel view renders over four scenes with each using three prompts. (Average of $4\times 3=12$ style transfer results.) Our method shows the best values for CLIP-TIDS, CLIP-DC, and user preference.
CLIP-TIDS $\uparrow$ CLIP-DC $\uparrow$ User Pref.
Instruct-NeRF2NeRF $0.081$ $0.871$ $15.3\%$
@@
5. Limitations and Future Work
While our method may apply artistic style transfer to various 3D scenes, including large-scale outdoor environments, there are several limitations to be considered. Depending on the strength of the stylization, there may be minor differences between stylized training images and NeRF renderings due to texture variations in the stylized multi-view images (See figure 6 (d) vs (f)). While a guidance scale of around 7.5-22.5 produces plausible results, a trade-off exists between stylization strength and view consistency. Thin structures such as plants and trees in the background or delicate texture patterns are also challenging to reconstruct due to ambiguities in the stylized multi-view images. For the same reason, our method will struggle to learn fine details if there is too much variation in the training images (e.g. different people or objects in the background, random patterns of clouds in the sky). As the style-aligned diffusion pipeline is conditioned on depth maps, significant editing of geometry is also difficult.
We think our approach is applicable to other types of 3D representations such as 3D Gaussian Splatting (Kerbl et al., 2023) and extendable to more features such as scene relighting and deformation, which are exciting directions for further exploration.
6. Conclusion
@@
Acknowledgements.
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015, and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. We want to thank Instruct-NeRF2NeRF (Haque et al., 2023) authors for sharing their dataset, and Yuki Kato, Shinji Terakawa, and Yoshiaki Tahara for assisting with data capturing.
@@ Appendix A Additional Implementation Details
A.1. NeRF Pipeline
We pre-train the ”nerfacto” model implemented in Nerfstudio (Tancik et al., 2023) for 60,000 iterations and then fine-tune for 15,000 iterations. We use the default ”nerfacto” losses; RGB pixel loss $\mathcal{L}_{rgb}$, distortion loss $\mathcal{L}_{dist}$ (Barron et al., 2022), interlevel loss $\mathcal{L}_{inter}$ (Barron et al., 2022), orientation loss $\mathcal{L}_{orien}$ (Verbin et al., 2022), and predicted normal loss $\mathcal{L}_{normal}$ (Verbin et al., 2022) for pre-training (equation 12). During fine-tuning, we disable the RGB pixel loss $\mathcal{L}_{rgb}$, orientation loss $\mathcal{L}_{orien}$, and predicted normal loss $\mathcal{L}_{normal}$ but add the Sliced Wasserstein Distance (SWD) loss (Heitz et al., 2021) (equation 13). The total loss function for each phase is as follows:
(12) $\mathcal{L}_{pre}=\mathcal{L}_{rgb}+\lambda_{1}\mathcal{L}_{dist}+\lambda_{2}\mathcal{L}_{inter}+\lambda_{3}\mathcal{L}_{orien}+\lambda_{4}\mathcal{L}_{normal}$
(13) $\mathcal{L}_{fine}=\mathcal{L}_{swd}+\lambda_{1}\mathcal{L}_{dist}+\lambda_{2}\mathcal{L}_{inter}$
where we use the default hyper-parameters $\lambda_{1}=0.002,\lambda_{2}=1.0,\lambda_{3}=0.0001,\lambda_{4}=0.001$ for most cases. A greater value for the distortion loss may work better if floating artifacts remain in the scene. We list a brief description of the ”nerfacto” losses.
  • Distortion Loss: The loss encourages the density along a ray to become compact, aiming to prevent floaters and background collapse. It was proposed in Mip-NeRF 360 (Barron et al., 2022).
  • Interlevel Loss: The loss allows the histograms of the point sampling proposal network and NeRF network to become more consistent. It was also proposed in Mip-NeRF 360 (Barron et al., 2022)
  • Orientation Loss: The loss aims to prevent ”foggy” surfaces by penalizing visible samples with predicted normals facing the ray direction. It was introduced in Ref-NeRF (Verbin et al., 2022).
@@ A.1. NeRF Pipeline
For detailed definitions of the ”nerfacto” losses, please see Mip-NeRF (Barron et al., 2022) and Ref-NeRF (Verbin et al., 2022).
We run all our experiments with Python 3.10 and CUDA 11.8 on a single NVIDIA H100. Large-scale outdoor 360 scenes (such as ”Campsite” or ”Farm,” consisting of 174 multi-view training images) takes -15 minutes for pre-training with -6GB GPU memory and fine-tuning takes -30 minutes with -20GB GPU memory (which depends on the number of ray samples).
The SWD loss $\mathcal{L}_{swd}$ is applied to $64\times 64$ patches sampled during NeRF fine-tuning. Although we found that $64\times 64$ empirically works sufficiently, one may change the patch size accordingly. While we optionally enable SDEdit (Meng et al., 2021) in our style-aligned diffusion pipeline, we recognized that depth maps are often enough for conditioning on the original views. In such cases, we use $1.0$ as the strength for SDEdit.
A.2. SWD Implementation
We follow the implementation of (Heitz et al., 2021) using the first 12 layers of VGG19 with uniformly sampled random projections for the SWD calculation.
Appendix B Additional Comparison of Loss Functions
We provide comparisons against two related loss functions: the Gram loss and Learned Perceptual Image Patch Similarity (LPIPS), and discuss the relative effectiveness of the SWD loss. Please see figure 9 for a visual comparison. We also show quantitative evaluation results in table 3.
Figure 9. Loss function comparison. We show novel view renders from fine-tuned NeRF models trained with different loss functions: (c) Ours (SWD Loss), (d) LPIPS Loss, and (e) Gram Loss. Gram loss introduces noticeable artifacts. While the LPIPS variant performs better than the Gram loss version, our NeRF render results are more similar to the stylized reference image with fewer artifacts.
Since style is known to be well represented by the feature maps of VGG19, we are interested in a loss term that accurately captures their distributions. Given the sets of feature distributions $p$ and $\hat{p}$ for the corresponding images $I$ and $\hat{I}$, the formal objective is to employ a loss term $\mathcal{L}$ satisfying equation 14 where $l$ denotes the layer number of VGG19. In short, we choose SWD loss (i.e., $\mathcal{L}=\mathcal{L}_{SW}$) as it can capture the complete stationary statistics of VGG19 feature distributions.
(14) $\mathcal{L}(I,\hat{I})=0\implies p^{l}=\hat{p}^{l},\forall l\in{1,\ldots,L}$
B.1. Gram Loss
Gram loss introduced by (Gatys et al., 2015) is defined as the mean-squared error between Gram matrices of the feature distributions:
(15) $\mathcal{L}_{Gram}(I,\hat{I})=\sum^{L}_{l}\frac{1}{N^{2}_{l}}\left\|G^{l}-\hat{G}^{l}\right\|^{2}$
where $G^{l}$ and $\hat{G}^{l}$ denote the Gram matrices of feature maps from images $I$ and $\hat{I}$ at layer $l$. Given a feature map of $M_{l}$ pixels with $N_{l}$ channels, an element $G_{i,j}$ of the Gram matrix $G^{l}\in\mathbb{R}^{N_{l}\times N_{l}}$ is defined as the second-order cross moment between features at channel $i$ and $j$.
(16) $G^{l}_{i,j}=\frac{1}{M_{l}}\sum^{M_{l}}_{m=1}F^{l}_{m}[i]F^{l}_{m}[j]$
Although Gram loss is often utilized as a convenient style loss for its capability to capture the feature statistics, Gram loss cannot capture the full distribution of features, resulting in some artifacts, whereas Wasserstein loss is able to capture the complete target distribution.
@@ B.1. Gram Loss
B.2. LPIPS
LPIPS (Zhang et al., 2018) is a metric developed for measuring perceptual similarity that well agrees with human evaluations. LPIPS is calculated with a pre-trained model that takes averaged feature maps as input and is trained via cross-entropy loss based on human-judged data. While LPIPS excels at capturing the perturbation- invariant perceptual similarity of patches, SWD better represents the raw VGG19 feature distributions, which makes it more appropriate for style transfer tasks. Replacing SWD with LPIPS produces mild artifacts.
Table 3. Additional Quantitative Comparison. We show CLIP-TIDS, CLIP-DC, and the averaged warping error (MSE) measured across rendered view frames from novel camera trajectories. The values are the average of two scenes using five prompts.
Gram Loss LPIPS Loss Ours (SWD Loss)
CLIP-TIDS $\uparrow$ $0.127$ $0.115$ $\mathbf{0.162}$
CLIP-DC $\uparrow$ 0.868 0.887 $\mathbf{0.928}$
Warp Error $\downarrow$ 0.323 0.332 $\mathbf{0.337}$
Appendix C CLIP Text-Image Direction Similarity
CLIP Text-Image Direction Similarity (CLIP TIDS) (Gal et al., 2022) is a metric for evaluating how well the change in the stylized image is aligned with the user-provided text prompt. Given the CLIP image and text encoder $E_{I},E_{T}$, CLIP TIDS between the source and the stylized image $I_{source},I_{style}$ is calculated as:
(18) $\Delta I=E_{I}(I_{style})-E_{I}(I_{source}),\Delta T=E_{T}(t_{style})-E_{T}(t_{source})$
[…]
where $t_{source},t_{style}$ are text descriptions describing the images (e.g. $t_{source}=$”A photo of a person”, $t_{style}=$”A person like Vincent Van Gogh”).
Appendix D Effect of Text Guidance Scale
In figure 10, we show an example of NeRF renderings using the same style prompt ”A person like Vincent Van Gogh” but different text guidance scales. We may see that style strength is controllable while keeping the original content structure grounded on the original image. We also show CLIP TIDS for each text guidance scale in table 4. A stronger text guidance scale will lead to higher CLIP TIDS.
Table 4. CLIP-TIDS Comparison. CLIP-TIDS values over test renders for each text guidance scale are shown. We can verify that a stronger text guidance scale results in higher CLIP TIDS values.
Text Guidance $s=7.5$ $s=15.0$ $s=22.5$ $s=30.0$ $s=37.5$
@@ Appendix D Effect of Text Guidance Scale
Figure 10. Effect of Text Guidance Scale. We show some NeRF rendering results for the prompt ”A person like Vincent Van Gogh” using various text guidance scales ($s=7.5,15.0,22.5,30.0,37.5$).
Figure 11. Limitations. Due to remaining ambiguities in the stylized multi-view images, fluctuating objects such as clouds may lose their detailed shape in the fine-tuned NeRF renderings. We wish to improve on this in our future work.
Appendix E Limitations
While our method can effectively perform overall style transfer to 3D scenes, it is still difficult to reconstruct a detailed structure of fluctuating objects within the stylized multi-view images. In figure 11, for example, (c) the clouds in the fine-tuned NeRF scene show a fractal-like pattern, which is different from (b) the clouds illustrated in the stylized image. This phenomenon is due to the ambiguities of cloud positions or shapes appearing in the stylized images generated by the style-aligned diffusion pipeline (Hertz et al., 2023b). We leave the development of a more robust content structure-preserving style transfer technique as future work.